# Multifractality of Many-Body Non-Hermitian Skin Effect

### v2

@@ Abstract
In this Letter, we elucidate that the skin effect manifests multifractality in non-Hermitian strongly correlated systems, thereby providing a distinctive hallmark of the many-body skin effect. We also show that the many-body skin effect can coexist with spectral statistics of random matrices. This contrasts with multifractality associated with the many-body localization, which is incompatible with ergodicity. In addition to non-Hermitian Hamiltonians, we demonstrate multifractality of the many-body skin effect within the Lindblad master equation. Our work reveals a defining characteristic of the many-body skin effect and underscores its fundamental role in open quantum many-body systems.
Multifractal scaling.—We consider a normalized wave function |$\ket{\psi}$ in a given computational basis |$\ket{n}$’s, |$\ket{\psi}=\sum_{n=1}^{\cal N}\psi_{n}\ket{n}$ ($\psi_{n}\in\mathbb{C}$), with the Hilbert space dimension $\mathcal{N}$. From the $q$th moments of this wave function, we introduce the $q$th participation entropy as
@@
For $q=2$, this reduces to the inverse participation ratio $S_{2}=-\ln\left(\sum_{n=1}^{\cal N}\left|\psi_{n}\right|^{4}\right)$.
Importantly, the participation entropy $S_{q}$ quantifies the localization properties in the given Hilbert space. For perfectly delocalized states, we have $S_{q}=\ln\mathcal{N}$. For states localized in a finite region of the Hilbert space, by contrast, $S_{q}$ no longer depends on $\mathcal{N}$. Between these two opposite regimes, states can exhibit the intermediate behavior $S_{q}=D_{q}\ln\mathcal{N}$ with $10<D_{q}<$ 1, implying that they are extended with nontrivial occupation in the Hilbert space—multifractality. We employ participation entropy, instead of entanglement entropy, to directly capture such localization properties. Here, $D_{q}\coloneqq S_{q}/\ln\mathcal{N}$ quantifies the effective dimension of the wave function occupying the Hilbert space.
@@
Model.—To capture a general feature of the many-body skin effect, we study the following nonintegrable non-Hermitian spin chain:
@@
For $J=g=h=0$, this model reduces to a non-Hermitian free fermionic model introduced by Hatano and Nelson [106, *Hatano-Nelson-97]. This model is the simplest model subject to the skin effect [51, 26, 33]. In fact, while all single-particle eigenstates form delocalized Bloch waves under the periodic boundary conditions, they are localized at either edge under the open boundary conditions. We find that these skin modes accompany the vanishing multifractal dimension $D_{q}=0$, indicating the perfect localization in the single-particle Hilbert space [108].
@@
Multifractality.—Through exact diagonalization, we calculate the complex spectrum and multifractal dimension $D_{q=2}$ for each right eigenstate under both periodic and open boundary conditions [Fig. 1 (a, b)]. While multifractality generally depends on the choice of the computational basis, we here consider the spin configuration. We choose the model parameters to ensure nonintegrability in the Hermitian limit [100] and investigate generic excited eigenstates instead of special eigenstates including the ground state. The complex spectrum undergoes substantial changes depending on the different boundary conditions. While many eigenstates exhibit large multifractal dimensions $D_{2}\simeq 1$ under the periodic boundary conditions, their counterparts under the open boundary conditions exhibit much smaller $D_{2}$, suggesting the skin effect. Multifractal dimensions $D_{2}$ quantify the degree of localization dependent on many-body eigenenergies $E$. Figure 1 (c, d) shows the distributions of multifractal dimensions $D_{2}$. Under both boundary conditions, $D_{2}$ realizes the maximum around the center of the many-body spectrum. Notably, $D_{2}$ deviates from unity even under the periodic boundary conditions, which should stem from the locality constraints in a similar manner to Hermitian quantum many-body systems [109, *Haque-22].
@@
As shown in Fig. 1 (a-d), the skin effect influences all many-body eigenstates. To capture this characteristic feature of the skin effect, we obtain the participation entropy $\braket{S_{q=2}}$ averaged over all right eigenstates as a function of the Hilbert space dimension $\mathcal{N}$ [Fig. 1 (e, f)] [111]. Fitting $\braket{S_{2}}$ by $\braket{S_{2}}=\braket{D_{2}}\ln\mathcal{N}+\braket{c_{2}}$, we obtain the dependence of the average multifractal dimension $\braket{D_{2}}$ and its subleading term $\braket{c_{2}}$ on $\gamma$ [Fig. 1 (g, h)]. Whereas $\braket{D_{2}}$ remains nearly constant under the periodic boundary conditions, it decreases for larger $\gamma$ under the open boundary conditions. This signifies the stronger skin effect in the many-body Hilbert space. Importantly, the many-body skin effect does not necessitate disorder, as opposed to the many-body localization. For small non-Hermiticity $\gamma\lesssim 0.2$, $\braket{D_{2}}$ seems insensitive to the boundary conditions, implying the absence of the skin effect.
@@
Several recent works studied local particle number distributions subject to the skin effect [70, 72, 74, 75, 76, 84, 85, 86, 87, 90, 91]. However, the localization of many-body skin modes should not be captured in real space but in many-body Hilbert space. The significant difference in multifractal dimensions between the different boundary conditions provides a quantitative measure of the skin effect inherent in non-Hermitian many-body systems. Additionally, we investigate multifractality in the interacting Hatano-Nelson model [85, 86, 112, 113, 114, 108]. Despite integrability, the many-body skin effect manifests itself as multifractality, akin to the nonintegrable model in Eq. (2). This further shows the generality of multifractality as a characteristic of the many-body skin effect.
@@
Spectral statistics.—Similarly to many-body skin modes, many-body-localized modes exhibit multifractality. However, we find a crucial distinction in quantum chaotic behavior, especially spectral statistics. Several recent works developed measures of chaos in open quantum systems [115, *Grobe-89, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128]. While the spectral statistics universally follow the random-matrix statistics in the chaotic regime, they instead follow the Poisson statistics in the integrable regime, providing a diagnosis of the dissipative quantum chaos.
@@
for an ordered set of singular values $s_{n}$’s ($n=1,2,\cdots,\mathcal{N}$[129]. Here, to break unwanted symmetry, we add a small disordered term $\sum_{i=1}^{L}\varepsilon_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}$ with a random number $\varepsilon_{i}$ distributed uniformly in $\left[-0.1,0.1\right]$ for each site $i$, which is expected not to affect multifractality significantly. Under both periodic and open boundary conditions, the singular-value statistics conform to the statistics of non-Hermitian random matrices [130], indicating the dissipative quantum chaos even in the presence of the skin effect. We further compute the statistics of complex level-spacing ratios, $z_{n}\coloneqq(E_{n}^{\rm NN}-E_{n})/(E_{n}^{\rm NNN}-E_{n})$ with the nearest-neighbor eigenvalue $E_{n}^{\rm NN}$ and next-to-nearest-neighbor eigenvalue $E_{n}^{\rm NNN}$ for each complex eigenvalue $E_{n}$ [123], conforming to the random-matrix statistics [Fig. 2 (b, c)] [131].
@@
Liouvillian skin effect.—Multifractality accompanies the many-body skin effect even within the quantum master equation. We investigate Markovian open quantum systems described by $d\rho/dt=\mathcal{L}\left(\rho\right)$ with the Lindbladian [132, 133, 134]
@@
We exactly diagonalize the Lindbladian $\mathcal{L}$ and obtain the complex spectrum and multifractal dimension $D_{q=2}$ for each right eigenoperator (Fig. 3). Depending on whether we impose the periodic or open boundary conditions, the complex spectrum differs substantially, signifying the Liouvillian skin effect. For both boundary conditions, the average multifractal dimensions $\braket{D_{2}}$ deviate from unity, likely due to the limited system size. Nevertheless, $\braket{D_{2}}$ under the open boundary conditions is significantly smaller than $\braket{D_{2}}$ under the periodic boundary conditions, demonstrating multifractality accompanied by the Liouvillian skin effect. Unlike the previous case for the non-Hermitian Hamiltonian in Eq. (2), $\braket{D_{2}}$ exhibits nonmonotonic dependence on the dissipation strength $\gamma$. This is because the Lindbladian dissipation incorporates both many-body interactions and nonreciprocity, competing with each other.
@@
Discussions.—Despite the considerable recent interest in the non-Hermitian skin effect, its many-body characterization has remained unestablished. In this Letter, we have uncovered a connection between two crucial but previously unrelated physical concepts—multifractality and skin effect. Multifractality provides a a defining feature of the skin effect in the many-body Hilbert space. In contrast to the many-body localization, multifractality due to the many-body skin effect can coexist with the random-matrix spectral statistics. This lets us revisit a fundamental relationship between multifractality and ergodicity, and reveals a unique role of the skin effect in open quantum systems.
@@
While we have focused on multifractality defined solely by right eigenstates in this Letter, we may also introduce multifractality by the combination of right and left eigenstates. For comparison, we also calculate multifractal dimensions from both right and left eigenoperators for the many-body Lindbladian [108], which may be relevant to the relaxation dynamics [94, 96]. It is worthwhile to further study such different measures of multifractality systematically. Moreover, recent years have seen various types of nonequilibrium phase transitions in open quantum systems. In our non-Hermitian spin model in Eq. (2), the average multifractal dimension $\braket{D_{2}}$ under the open boundary conditions remains at unity for $\gamma\lesssim 0.2$ but deviates from unity for $\gamma\gtrsim 0.2$, implying a possible phase transition induced by the interplay between non-Hermiticity and many-body interactions. It merits further study to determine whether this is a sharp phase transition or a crossover in the infinite-size limit.
@@ Acknowledgements.
We thank Ryusuke Hamazaki, Takashi Mori, Shoki Sugimoto, and Tsuneya Yoshida for helpful discussions. We appreciate the long-term workshop “Quantum Information, Quantum Matter and Quantum Gravity” (YITP-T-23-01) held at Yukawa Institute for Theoretical Physics (YITP), Kyoto University. S.H. acknowledges travel support from MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A through the “Extreme Universe” collaboration (Grant Nos. 21H05182 and 22H05247). S.H. is supported by JSPS Research Fellow, JSPS Overseas Challenge Program for Young Researchers, and MEXT WISE Program. K.K. is supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” No. 24H00945.
Supplemental Material for “Multifractality of Many-Body Non-Hermitian Skin Effect”
I Noninteracting Hatano-Nelson model
@@
$H=2t\sum_{i=1}^{L}\left[\left(1+\gamma\right)c_{i+1}^{{\dagger}}c_{i}+\left(1-\gamma\right)c_{i}^{{\dagger}}c_{i+1}\right],$ (I.1)
@@
$H=\frac{t}{2}\sum_{i=1}^{L}\left[\left(1+\gamma\right)\sigma_{i}^{-}\sigma_{i+1}^{+}+\left(1-\gamma\right)\sigma_{i}^{+}\sigma_{i+1}^{-}\right].$ (I.2)
@@
|$\ket{\psi_{n}}=\frac{1}{\sqrt{L}}\sum_{i=1}^{L}e^{\text{i}k_{n}i}c_{i}^{{\dagger}}\ket{\rm vac},\quad k_{n}\coloneqq\frac{2\pi n}{L}\quad\left(n=0,1,2,\cdots,L-1\right)$ (I.3)
with the vacuum |$\ket{\rm vac}$ of fermions (i.e., $c_{i}\ket{\rm vac}=0$ for all $i$). The $q$th participation entropy in the single-particle Hilbert space is obtained as
$S_{q}=\frac{1}{1-q}\ln\left(\sum_{i=1}^{L}\left|\frac{e^{\text{i}k_{n}i}}{\sqrt{L}}\right|^{2q}\right)=\ln L$ (I.4)
@@
|$\ket{\psi_{n}}\propto\sum_{i=1}^{L}\left(\beta^{i}\sin\left(k_{n}i\right)\right)c_{i}^{{\dagger}}\ket{\rm vac},\quad\beta\coloneqq\sqrt{\frac{1+\gamma}{1-\gamma}},\quad k_{n}\coloneqq\frac{\pi n}{L+1}\quad\left(n=1,2,\cdots,L\right).$ (I.5)
All of these eigenstates are localized at the right (left) edge for $\gamma>0$ ($\gamma<0$), which is a signature of the non-Hermitian skin effect in the single-particle Hilbert space. For simplicity, we approximate these single-particle eigenstates as
|$\ket{\psi_{n}}\simeq\frac{1}{\sqrt{N}}\sum_{i=1}^{L}\left(\beta e^{\text{i}k_{n}}\right)^{i}c_{i}^{{\dagger}}\ket{\rm vac}$ (I.6)
@@
$N=\sum_{i=1}^{L}\beta^{2i}=\frac{\beta^{2}\left(\beta^{2L}-1\right)}{\beta^{2}-1}.$ (I.7)
@@
$S_{q}\simeq\frac{1}{1-q}\ln\left(\sum_{i=1}^{L}\left|\frac{\left(\beta e^{\text{i}k_{n}}\right)^{i}}{\sqrt{N}}\right|^{2q}\right)=\frac{1}{1-q}\ln\left(\left(\frac{\beta^{2}-1}{\beta^{2}\left(\beta^{2L}-1\right)}\right)^{q}\frac{\beta^{2q}\left(\beta^{2qL}-1\right)}{\beta^{2q}-1}\right)\simeq\frac{1}{1-q}\ln\left(\frac{\left(\beta^{2}-1\right)^{q}}{\beta^{2q}-1}\right).$ (I.8)

### v3

@@ I Introduction
In this work, we elucidate that the skin effect manifests multifractality in non-Hermitian strongly correlated systems, thereby providing a distinctive hallmark of the many-body skin effect. We also show that the many-body skin effect can coexist with spectral statistics of random matrices. This contrasts with multifractality associated with the many-body localization, which is incompatible with ergodicity. In addition to non-Hermitian Hamiltonians, we demonstrate multifractality of the many-body skin effect within the Lindblad master equation. Our work reveals a defining characteristic of the many-body skin effect and underscores its fundamental role in open quantum many-body systems.
II Multifractal scaling
We consider a normalized wave function |$\ket{\psi}$ in a given computational basis |$\ket{n}$’s,
|$\ket{\psi}=\sum_{n=1}^{\cal N}\psi_{n}\ket{n}\quad(\psi_{n}\in\mathbb{C}),$ (1)
with the Hilbert space dimension $\mathcal{N}$. From the $q$th moments of this wave function, we introduce the $q$th participation entropy as
@@ II Multifractal scaling
For $q=2$, this reduces to the inverse participation ratio
$S_{2}=-\ln\left(\sum_{n=1}^{\cal N}\left|\psi_{n}\right|^{4}\right).$ (3)
Importantly, the participation entropy $S_{q}$ quantifies the localization properties in the given Hilbert space. For perfectly delocalized states, we have $S_{q}=\ln\mathcal{N}$. For states localized in a finite region of the Hilbert space, by contrast, $S_{q}$ no longer depends on $\mathcal{N}$. Between these two opposite regimes, states can exhibit the intermediate behavior
$S_{q}=D_{q}\ln\mathcal{N}\quad\left(0<D_{q}<1\right),$ (4)
implying that they are extended with nontrivial occupation in the Hilbert space—multifractality. We employ participation entropy, instead of entanglement entropy, to directly capture such localization properties. Here, $D_{q}\coloneqq S_{q}/\ln\mathcal{N}$ quantifies the effective dimension of the wave function occupying the Hilbert space.
@@
III Model
To capture a general feature of the many-body skin effect, we study the following nonintegrable non-Hermitian spin chain:
@@ III Model
For $J=g=h=0$, this model reduces to a non-Hermitian free fermionic model introduced by Hatano and Nelson [106, *Hatano-Nelson-97]. This model is the simplest model subject to the skin effect [51, 26, 33]. In fact, while all single-particle eigenstates form delocalized Bloch waves under the periodic boundary conditions, they are localized at either edge under the open boundary conditions. In Appendix A, we show that these skin modes accompany the vanishing multifractal dimension $D_{q}=0$, indicating the perfect localization in the single-particle Hilbert space.
IV Multifractality
@@
Through exact diagonalization, we calculate the complex spectrum and multifractal dimension $D_{q=2}$ for each right eigenstate under both periodic and open boundary conditions [Fig. 1 (a, b)]. While multifractality generally depends on the choice of the computational basis, we here consider the spin configuration. We choose the model parameters to ensure nonintegrability in the Hermitian limit [100] and investigate generic excited eigenstates instead of special eigenstates including the ground state. The complex spectrum undergoes substantial changes depending on the different boundary conditions. While many eigenstates exhibit large multifractal dimensions $D_{2}\simeq 1$ under the periodic boundary conditions, their counterparts under the open boundary conditions exhibit much smaller $D_{2}$, suggesting the skin effect. Multifractal dimensions $D_{2}$ quantify the degree of localization dependent on many-body eigenenergies $E$. Figure 1 (c, d) shows the distributions of multifractal dimensions $D_{2}$. Under both boundary conditions, $D_{2}$ realizes the maximum around the center of the many-body spectrum. Notably, $D_{2}$ deviates from unity even under the periodic boundary conditions, which should stem from the locality constraints in a similar manner to Hermitian quantum many-body systems [108, *Haque-22].
@@
As shown in Fig. 1 (a-d), the skin effect influences all many-body eigenstates. To capture this characteristic feature of the skin effect, we obtain the participation entropy $\braket{S_{q=2}}$ averaged over all right eigenstates as a function of the Hilbert space dimension $\mathcal{N}$ [Fig. 1 (e, f)]. The multifractal dimension averaged exclusively over midspectrum eigenstates shows no significant deviations. Fitting $\braket{S_{2}}$ by
$\braket{S_{2}}=\braket{D_{2}}\ln\mathcal{N}+\braket{c_{2}},$ (6)
we obtain the dependence of the average multifractal dimension $\braket{D_{2}}$ and its subleading term $\braket{c_{2}}$ on $\gamma$ [Fig. 1 (g, h)]. Whereas $\braket{D_{2}}$ remains nearly constant under the periodic boundary conditions, it decreases for larger $\gamma$ under the open boundary conditions. This signifies the stronger skin effect in the many-body Hilbert space. Importantly, the many-body skin effect does not necessitate disorder, as opposed to the many-body localization. For small non-Hermiticity $\gamma\lesssim 0.2$, $\braket{D_{2}}$ seems insensitive to the boundary conditions, implying the absence of the skin effect. In Appendix B, we provide additional numerical results.
@@
Several recent works studied local particle number distributions subject to the skin effect [70, 72, 74, 75, 76, 84, 85, 86, 87, 90, 91]. However, the localization of many-body skin modes should not be captured in real space but in many-body Hilbert space. The significant difference in multifractal dimensions between the different boundary conditions provides a quantitative measure of the skin effect inherent in non-Hermitian many-body systems. Additionally, in Appendix C, we investigate multifractality in the interacting Hatano-Nelson model [85, 86, 110, 111, 112]. Despite integrability, the many-body skin effect manifests itself as multifractality, akin to the nonintegrable model in Eq. (5). This further shows the generality of multifractality as a characteristic of the many-body skin effect.
V Spectral statistics
@@
Similarly to many-body skin modes, many-body-localized modes exhibit multifractality. However, we find a crucial distinction in quantum chaotic behavior, especially spectral statistics. Several recent works developed measures of chaos in open quantum systems [113, *Grobe-89, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126]. While the spectral statistics universally follow the random-matrix statistics in the chaotic regime, they instead follow the Poisson statistics in the integrable regime, providing a diagnosis of the dissipative quantum chaos.
@@
for an ordered set of singular values $s_{n}$’s ($n=1,2,\cdots,\mathcal{N}$[127]. Here, to break unwanted symmetry, we add a small disordered term $\sum_{i=1}^{L}\varepsilon_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}$ with a random number $\varepsilon_{i}$ distributed uniformly in $\left[-0.1,0.1\right]$ for each site $i$, which is expected not to affect multifractality significantly. Under both periodic and open boundary conditions, the singular-value statistics conform to the statistics of non-Hermitian random matrices, indicating the dissipative quantum chaos even in the presence of the skin effect.
We further compute the statistics of complex level-spacing ratios [121],
$z_{n}\coloneqq\frac{E_{n}^{\rm NN}-E_{n}}{E_{n}^{\rm NNN}-E_{n}}$ (8)
where $E_{n}^{\rm NN}$ denotes the nearest-neighbor eigenvalue, and $E_{n}^{\rm NNN}$ the next-to-nearest-neighbor eigenvalue for each complex eigenvalue $E_{n}$. The numerical results conform to the random-matrix statistics, as shown in Fig. 2 (b, c). The slight deviation for the periodic boundary conditions should be due to average translation symmetry. We note that the non-Hermitian Hamiltonian in Eq. (5) respects time-reversal symmetry (i.e., $H^{*}=H$). Consequently, its spectral statistics follow the random-matrix statistics for class AI [52, 119]. Time-reversal symmetry also requires the complex spectrum to be symmetric about the real axis [see Fig. 1 (a, b)].
@@
VI Liouvillian skin effect
Multifractality accompanies the many-body skin effect even within the quantum master equation. We investigate Markovian open quantum systems described by $d\rho/dt=\mathcal{L}\left(\rho\right)$ with the Lindbladian [128, 129, 130]
@@ VI Liouvillian skin effect
We exactly diagonalize the Lindbladian $\mathcal{L}$ and obtain the complex spectrum and multifractal dimension $D_{q=2}$ for each right eigenoperator, as shown in Fig. 3. Depending on whether we impose the periodic or open boundary conditions, the complex spectrum differs substantially, signifying the Liouvillian skin effect. For both boundary conditions, the average multifractal dimensions $\braket{D_{2}}$ deviate from unity, likely due to the limited system size. Nevertheless, $\braket{D_{2}}$ under the open boundary conditions is significantly smaller than $\braket{D_{2}}$ under the periodic boundary conditions, demonstrating multifractality accompanied by the Liouvillian skin effect. Unlike the previous case for the non-Hermitian Hamiltonian in Eq. (5), $\braket{D_{2}}$ exhibits nonmonotonic dependence on the dissipation strength $\gamma$. This is because the Lindbladian dissipation incorporates both many-body interactions and nonreciprocity, competing with each other.
Figure 4: Multifractality of the Liouvillian skin effect ($t=1/\sqrt{2}$, $J=\left(5-\sqrt{3}\right)/3$, $h=\left(1+\sqrt{5}\right)/4$). In both bra and ket spaces, the particle number is chosen as $L/2$ and $\left(L-1\right)/2$ for even and odd $L$, respectively. Participation entropy $\braket{S_{2}^{\rm RL}}$ averaged over all right and left eigenoperators as a function of the double Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=5,6,\cdots,9$).
Moreover, we introduce participation entropy and multifractal dimensions based on the combination of right and left eigenoperators, instead of those based solely on right eigenoperators. We consider right and left eigenoperators of non-Hermitian superoperators,
|$\ket{r_{\alpha}}=\sum_{n=1}^{\cal N}r_{n}\ket{n},\quad\ket{l_{\alpha}}=\sum_{n=1}^{\cal N}l_{n}\ket{n}\quad\left(r_{n},l_{n}\in\mathbb{C}\right),$ (12)
$\braket{r_{\alpha}}{r_{\alpha}}=\braket{l_{\alpha}}{l_{\alpha}}=1.$ (13)
$S_{q}^{\rm RL}\coloneqq\frac{1}{1-q}\ln\left(\sum_{n=1}^{\cal N}\left|r_{n}l_{n}\right|^{q}\right),$ (14)
and the multifractal dimension as
$D_{q}^{\rm RL}\coloneqq\frac{S_{q}^{\rm RL}}{\ln\mathcal{N}}=\frac{1}{1-q}\frac{1}{\ln\mathcal{N}}\ln\left(\sum_{n=1}^{\cal N}\left|r_{n}l_{n}\right|^{q}\right).$ (15)
We calculate the participation entropy $\braket{S_{q=2}^{\rm RL}}$ averaged over all right and left eigenoperators (Fig. 4). As we increase the double Hilbert space dimension $\mathcal{N}$, the average participation entropy $\braket{S_{2}^{\rm RL}}$ under the periodic boundary conditions grows more slowly than $\braket{S_{2}^{\rm RL}}$ under the open boundary conditions. This behavior is qualitatively similar to the behavior of the average participation entropy $\braket{S_{2}}$ defined solely by right eigenoperators (see Fig. 3). Additionally, as we increase the dissipation strength $\gamma$, the average participation entropy $\braket{S_{2}^{\rm RL}}$ remains almost the same for the periodic boundary conditions but increases for the open boundary conditions. It is worthwhile to further study the multifractal dimension $D_{q}^{\rm RL}$ defined by both right and left eigenoperators in a more systematic manner.
VII Discussions
@@
Despite the considerable recent interest in the non-Hermitian skin effect, its many-body characterization has remained unestablished. In this work, we have uncovered a connection between two crucial but previously unrelated physical concepts—multifractality and skin effect. Multifractality provides a defining feature of the skin effect in the many-body Hilbert space. In contrast to the many-body localization, multifractality due to the many-body skin effect can coexist with the random-matrix spectral statistics. This lets us revisit a fundamental relationship between multifractality and ergodicity, and reveals a unique role of the skin effect in open quantum systems.
@@
While we have mainly focused on multifractality defined solely by right eigenstates in this work, we also introduce multifractality by the combination of right and left eigenstates in Sec. VI, which may be relevant to the relaxation dynamics [94, 96]. It is worthwhile to further study such different measures of multifractality systematically. Moreover, recent years have seen various types of nonequilibrium phase transitions in open quantum systems. In our non-Hermitian spin model in Eq. (5), the average multifractal dimension $\braket{D_{2}}$ under the open boundary conditions remains at unity for $\gamma\lesssim 0.2$ but deviates from unity for $\gamma\gtrsim 0.2$, implying a possible phase transition induced by the interplay between non-Hermiticity and many-body interactions. It merits further study to determine whether this is a sharp phase transition or a crossover in the infinite-size limit.
@@ Acknowledgements.
We thank Ryusuke Hamazaki, Takashi Mori, Shoki Sugimoto, and Tsuneya Yoshida for helpful discussions. We appreciate the long-term workshop “Quantum Information, Quantum Matter and Quantum Gravity” (YITP-T-23-01) held at Yukawa Institute for Theoretical Physics (YITP), Kyoto University. S.H. acknowledges travel support from MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A through the “Extreme Universe” collaboration (Grant Nos. 21H05182 and 22H05247). S.H. is supported by JSPS Research Fellow No. 24KJ1445, JSPS Overseas Challenge Program for Young Researchers, and MEXT WISE Program. K.K. is supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” No. 24H00945.
Appendix A Noninteracting Hatano-Nelson model
@@
$H=2t\sum_{i=1}^{L}\left[\left(1+\gamma\right)c_{i+1}^{{\dagger}}c_{i}+\left(1-\gamma\right)c_{i}^{{\dagger}}c_{i+1}\right],$ (16)
@@
$H=\frac{t}{2}\sum_{i=1}^{L}\left[\left(1+\gamma\right)\sigma_{i}^{-}\sigma_{i+1}^{+}+\left(1-\gamma\right)\sigma_{i}^{+}\sigma_{i+1}^{-}\right].$ (17)
@@
|$\ket{\psi_{n}}=\frac{1}{\sqrt{L}}\sum_{i=1}^{L}e^{\text{i}k_{n}i}c_{i}^{{\dagger}}\ket{\rm vac},$ (18)
with $k_{n}\coloneqq 2\pi n/L$ ($n=0,1,2,\cdots,L-1$), and the vacuum |$\ket{\rm vac}$ of fermions (i.e., $c_{i}\ket{\rm vac}=0$ for all $i$). The $q$th participation entropy in the single-particle Hilbert space is obtained as
$S_{q}=\frac{1}{1-q}\ln\left(\sum_{i=1}^{L}\left|\frac{e^{\text{i}k_{n}i}}{\sqrt{L}}\right|^{2q}\right)=\ln L$ (19)
@@
|$\ket{\psi_{n}}\propto\sum_{i=1}^{L}\left(\beta^{i}\sin\left(k_{n}i\right)\right)c_{i}^{{\dagger}}\ket{\rm vac},\quad\beta\coloneqq\sqrt{\frac{1+\gamma}{1-\gamma}},$ (20)
with $k_{n}=\pi n/\left(L+1\right)$ ($n=1,2,\cdots,L$). All of these eigenstates are localized at the right (left) edge for $\gamma>0$ ($\gamma<0$), which is a signature of the non-Hermitian skin effect in the single-particle Hilbert space. For simplicity, we approximate these single-particle eigenstates as
|$\ket{\psi_{n}}\simeq\frac{1}{\sqrt{N}}\sum_{i=1}^{L}\left(\beta e^{\text{i}k_{n}}\right)^{i}c_{i}^{{\dagger}}\ket{\rm vac}$ (21)
@@
$N=\sum_{i=1}^{L}\beta^{2i}=\frac{\beta^{2}\left(\beta^{2L}-1\right)}{\beta^{2}-1}.$ (22)
@@
$S_{q}$ $\simeq\frac{1}{1-q}\ln\left(\sum_{i=1}^{L}\left|\frac{\left(\beta e^{\text{i}k_{n}}\right)^{i}}{\sqrt{N}}\right|^{2q}\right)$
$=\frac{1}{1-q}\ln\left(\left(\frac{\beta^{2}-1}{\beta^{2}\left(\beta^{2L}-1\right)}\right)^{q}\frac{\beta^{2q}\left(\beta^{2qL}-1\right)}{\beta^{2q}-1}\right)$
$\simeq\frac{1}{1-q}\ln\left(\frac{\left(\beta^{2}-1\right)^{q}}{\beta^{2q}-1}\right).$ (23)